Cord A is 6.00 m long, has breaking strength of 235. N, and is used to connect two balls of mass 500. g each. Cord B is identical to Cord A and has one end connected to one of the balls. The other end is held in order to swing the balls in a horizontal circle. As the balls are swung faster and faster,

(a) which of the cords will break first?

(b) At what angular speed, in revolutions per second, will the break occur?


1
Expert's answer
2021-04-13T06:40:57-0400

a) Cord A will break first.


b)


"F=m\omega^2R\\235=0.5(2\cdot 6)\omega^2\\\omega=6.26\frac{rad}{s}=0.996\frac{rev}{s}"